1.28 As a principal in the consulting firm where you have worked for 20 years, you have accumulated 5000 shares of company stock. One year ago, each share of stock was worth $40. The company has offered to buy back your shares for $225,000. At what interest rate would the firms offer be equiva- lent to the worth of the stock last year? 1.35
1.30 University tuition and fees can be paid by using one of two plans. Early-bird: Pay total amount due I year in advance and get a 10% discount. On-time: Pay total amount due when classes start. The cost of tuition and fees is $10,000 per year. (a) How much is paid in the early-bird plan? (b) What is the equivalent amount of the savings compared to the on-time payment at the time that the on-time payment is made?
Answer #1

Following are the answer:-

= 5000(40)(1 + i) = 225,000

1 + i = 1.125

i = 0.125

= 12.5% per year.

Same answer in other format , Following are the answer:-